# Data-driven tool wear prediction in milling, based on a process-integrated single-sensor approach

### v2

@@ IV-A Setup
For this experiment, the tool life cycle of a solid carbide square end mill with four teeth and a diameter of 12 mm is analyzed on two different machines: the first machine is the ’CHIRON FZ 15 S’ and the second machine the ’DMU 60 FD duoBLOCK’ . The workpiece material is hardened steel (42CrMo4, +QT), chosen for its demanding properties. The toolholder and sensor used for the experiment is the ’iTENDO²’ from Schunk, which integrates a sensor directly into the center of the toolholder housing (Fig. 4a). The sensor used is a single-axis MEMS acceleration sensor that measures radial acceleration ($a_{r}$). As the sensor is placed on the rotation axis of the toolholder, it is able to effectively measure the acceleration on the x-axis ($a_{x}$) and y-axis ($a_{y}$) over time, as it rotates with the toolholder and spindle speed. Fig. 4b illustrates the measurement behavior, which presents a top view of the iTENDO² in the xy plane with the axis of rotation being the z-axis.
@@
If the sensor is installed optimally, it does not measure tangential acceleration $a_{t}$. As a result, the measured accelerations are dynamically mixed depending on the angle of rotation, enriching the signal with more detailed acceleration information. However, this also makes the signal more challenging to interpret. The reason for this is that the rotation matrix multiplies frequencies in the time domain, as both $a_{x}$ and $a_{y}$ also consist of frequency components. In accordance with the convolution theorem, the frequencies in the frequency domain are also convolved. This particular sensor/toolholder combination is still chosen because the sensor’s position allows it to capture undamped process vibrations more effectively. In addition, the setup is easily transferable to other machines and processes, requiring only the replacement of the toolholder to record training data under similar conditions. This modularity aligns seamlessly with the key contribution of this research: the development of a novel TCM pipeline that enables models to be swiftly adapted to a range of industrial milling processes for wear classification.
@@ IV-B Design of Experiment
In this experiment, a flat surface is created on a metal block using side milling. This milling process is repeated continuously until the cutting tool shows significant wear. The milling operation is a down-milling approach, with coolant lubricant applied throughout the process. Milling is performed exclusively along the xy plane. Detailed process parameters are given in Table II and the visualization of the experiment is shown in Fig. 5.
@@
In order to accurately characterize the acceleration data, the flank wear of the end mill is measured at regular intervals. When the flank wear reached 0.2 mm, the tool is considered to be highly worn, indicating the completion of one tool life cycle and the process run restarts with a new tool. In addition to direct measurement, the machine operator provides a subjective label based on their experience. This label is the tool state, indicating when the tool wear had reached a level where the negative impact on the process becomes too great. The tool state is used for wear prediction in this research.
@@ V Evaluation
It can be seen that there are only small differences between the training data of 80% and 60%. This indicates that a good prediction of machine wear can be achieved by training on 60% training data, which corresponds to three tool life cycles. In addition, we see that the first group, using the STFT as input, performs better. However, due to the larger number of parameters and the deeper architecture, these models have a higher computational cost and memory requirement. Although the prediction accuracy or binary accuracy reflects the overall quality of the model, in reality it is more important how the model predicts wear over time, in particular when the model changes its prediction from not worn to worn. Fig. 7 shows the prediction of wear over time. The y-axis illustrates the categories of the binary classification, which are defined as follows:
$C:\{0,1\}\to\{\text{Not worn},\text{Worn}\}$ (9)
The wear subjectively labeled by the operator (true label) is shown in blue and the model prediction is shown in green. Depending on the plot, several tool life cycles are shown concatenated in one plot. When the label changes from Worn to Not worn, a new tool life cycle begins. For post-processing, a simple filter is applied to the classification results to address the logical inconsistency that a tool cannot briefly appear worn and then revert to an unworn state. This filter eliminates a small proportion of such outliers from the plots, thereby enhancing both the accuracy over time and the readability of the visual representations. However, prominent outliers remain visible in the plots. Fig. 7 illustrates that time accuracy improves as the size of the training set increases. It is noteworthy that certain models, such as ”ConvNeXt”, demonstrate relatively good time accuracy even with only 40% of the training data. The ”SVC”, despite its simplicity, achieves only slightly lower time accuracy compared to more advanced models. The results together demonstrate that a minimum of two, preferably three, tool life cycles is required to train a model with the capacity to make reliable predictions. This assumption is based on the premise that the transition from no wear to worn occurs with minimal time deviations. In practice, recognizing wear too late is undesirable. To mitigate this, the transition from ”not worn” to ”worn” could be adjusted slightly earlier in the labeling process, thereby introducing a calculated bias for the early prediction of wear. Alternatively, if avoiding such a bias is preferred and high time accuracy is still desired, training on data from four or more tool life cycles becomes necessary.
@@
It should be noted that the training data of the different splits are identical to those used in the heatmap in Fig. 6. It is evident that the models on the unseen machine exhibit reduced performance, yet the ConvNeXt model continues to demonstrate the most optimal outcomes. Interestingly, models trained with the largest dataset (80% training data) perform worse than those trained with smaller datasets. This suggests that the change of a machine has fundamentally altered the wear behavior, causing models trained on extensive data to overfit to the specific characteristics of the original machine. Models trained with 40% of the training data achieve the highest accuracy. An examination of the accuracy over time (Fig. 9) reveals that the models are capable of detecting changes in the DMU60FD dataset and successfully predict wear in some tool life cycles. Nevertheless, when considering all models collectively, while each tool life cycle is successfully recognized by at least one model, no single model achieves consistent accuracy across all cycles. This suggests that the findings derived from the training dataset can be applied to the DMU60FD dataset. However, no overarching patterns are identified that apply to all tool life cycles. In summary, the fundamental issue associated with TCM’s is also clear in this study, namely that the complex characteristics of each machine make it difficult to implement one prediction model across other unseen machines. To improve predictions, the models require additional training data from the specific machine.
@@ VI Conclusion
This paper addresses the question of how to develop a TCM system that is transferable to industrial applications. For this, a modular approach and a non-invasive setup is employed so that a TCM can be installed for different milling processes with minimal effort. For a non-invasive setup, a 1D acceleration sensor is used, and for modality the sensor is integrated into a toolholder. In this research, the acceleration signal is processed in both the time domain and the frequency domain, to generate valuable data input for wear prediction or TCM. The best performing TCM approach uses short-time Fourier transform as input and an automatic feature extraction through CNN architecture to predict wear. Given that the development of a TCM is more cost-effective when fewer training data are required, the study also investigates the amount of training data needed from a milling process to get a sufficient accuracy: To achieve the highest degree of accuracy in wear prediction, at least four tool life cycles should be included in the training data, two or three cycles for an acceptable accuracy. Among the models tested, the ConvNeXt model delivered the best results with 99% accuracy. Additionally, the models capacity to predict wear for an untrained milling machine is evaluated. Although models demonstrate some ability to transfer information and predict wear for certain tools, they struggle to fully capture the inherent complexity and variability of new milling machines, resulting in inconsistent wear predictions. Transfer learning will therefore be necessary for new processes. A potential next step is to investigate the quantity of data required for effective transfer learning.
@@ Appendix A Tables of this research
TABLE I: Feature names and their corresponding formulas
@@
Mean $\frac{1}{N}\sum_{i=1}^{N}x_{i}$
RMS (Root Mean Square) $\sqrt{\frac{1}{N}\sum_{i=1}^{N}x_{i}^{2}}$
SD (Standard Deviation) $\sqrt{\frac{1}{N-1}\sum_{i=1}^{N}(x_{i}-\bar{x})^{2}}$
Crest Factor $\frac{\max|x|}{\text{RMS}}$
Kurtosis $\frac{1}{N}\sum_{i=1}^{N}\left(\frac{x_{i}-\bar{x}}{\text{SD}}\right)^{4}$
Skewness $\frac{1}{N}\sum_{i=1}^{N}\left(\frac{x_{i}-\bar{x}}{\text{SD}}\right)^{3}$
Statistical Mode Most frequent value in the dataset
SD of Statistical Mode Standard deviation of the statistical mode
P2P (Peak-to-Peak) $\max(x)-\min(x)$
Energy of Power Spectral Density (PSD) $\frac{\sum{|X(f)|^{2}}}{N}$, where $X(f)$ is the PSD of $x(t)$
Power Spectral Entropy $-\sum_{n=1}^{N}P(n)\log_{2}P(n)$, where $P(n)$ is the probability distribution of a power spectrum
Relative Energy of aperiodic frequencies $\frac{\text{Energy}_{ap}}{\text{Energy}_{all}}$, where $ap$ is energy of the aperiodic frequencies
Auto-correlation $\frac{\sum_{i=1}^{N-\tau}(x_{i}-\bar{x})(x_{i+\tau}-\bar{x})}{\sum_{i=1}^{N}(x_{i}-\bar{x})^{2}}$
Higuchi Fractal Dimension Linear regression on the points $\{(log(1/k),log(L(k))\}$ to estimate the slope, where $k$ is the interval length or scale parameter and $L(k)$ the curve length
TABLE II: Process parameters from the measured acceleration data for the models

### v3

@@ IV-A Setup
Typical components used in a standard milling process were selected. For this experiment, the tool life cycle of a solid carbide square end mill with four flutes and a diameter of 12 mm is analyzed on two different machines: the first machine is the CHIRON FZ 15 S and the second machine is the DMU 60 FD duoBLOCK. The workpiece material is hardened steel (42CrMo4, +QT), chosen for its demanding properties.
As discussed previously, an accelerometer was selected due to its excellent balance of cost efficiency, ease of integration, and high-quality information for tool wear monitoring. To further enhance industrial application, this study employs a single-sensor configuration to minimize costs while maximizing integrability. However, using only one sensor introduces challenges that must be addressed through optimal sensor selection. Since accelerometers excel in capturing frequency-domain information, several key factors must be considered to ensure reliable wear detection:
  • High sampling rate to accurately capture rapid signal variations.
  • Broad frequency range and high frequency resolution to detect relevant vibration signatures.
  • High resonance frequency to avoid distortions and ensure accurate measurements.
  • Extremely low noise density to maintain signal clarity.
  • Positioned close to the cutting process to maximize sensitivity to process-related vibrations while minimizing external noise interference.
To meet these requirements, a tool holder-integrated sensor was selected, allowing direct measurement as close to the cutting process as possible while simplifying installation. Following a thorough evaluation, the iTENDO² from Schunk was identified as the optimal solution (Fig. 5a). The sensor used is a single-axis, low noise, high frequency MEMS acceleration sensor that measures radial acceleration ($a_{r}$). As the sensor is placed on the rotation axis of the toolholder, it is able to effectively measure the acceleration on the x-axis ($a_{x}$) and y-axis ($a_{y}$) over time, as it rotates with the toolholder and spindle speed. Fig. 5b illustrates the measurement behavior, which presents a top view of the iTENDO² in the xy plane with the axis of rotation being the z-axis.
@@
If the sensor is installed optimally, it does not measure tangential acceleration $a_{t}$. As a result, the measured accelerations are dynamically mixed depending on the angle of rotation, enriching the signal with more detailed acceleration information. However, this also makes the signal more challenging to interpret. The reason for this is that the rotation matrix multiplies frequencies in the time domain, as both $a_{x}$ and $a_{y}$ also consist of frequency components. In accordance with the convolution theorem, the frequencies in the frequency domain are also convolved. This particular sensor/toolholder combination is still chosen because the sensor’s position allows it to capture undamped process vibrations more effectively. In addition, the setup can be easily transferred to other machines and processes, requiring only replacement of the toolholder to record training data under similar conditions. This modularity aligns seamlessly with the key contribution of this research: the development of a novel TCM pipeline that enables models to be swiftly adapted to a range of industrial milling processes for wear classification.
@@ IV-B Design of Experiment
In this experiment, a flat surface is created on a metal block using side milling. This milling process is repeated continuously until the cutting tool shows significant wear. The milling operation is a down-milling approach, with coolant lubricant applied throughout the process. Milling is performed exclusively along the xy plane. Detailed process parameters are given in Tab. III and the visualization of the experiment is shown in Fig. 6.
@@
In order to accurately characterize the acceleration data, the flank wear of the end mill is measured at regular intervals. When the flank wear reached 0.2 mm, the tool is considered to be highly worn, indicating the completion of one tool life cycle and the process run restarts with a new tool. Flank wear is a reliable indicator of tool condition, directly influencing cutting performance. However, it is important to note that flank wear is just one of several wear mechanisms and its reliability as a universal wear metric is dependent on the specific milling conditions. This variability in performance makes it impossible to define a strict threshold for tool replacement based solely on flank wear measurements. Instead, an approximation is used to determine wear severity. Since this study aims to develop an industry-oriented solution, the model is not designed to predict flank wear itself but to classify the tool’s condition as either usable or worn. To enhance the labeling process, an experienced machine operator manually evaluates the tool’s condition and assigns a ToolState label, as illustrated in Fig. 7.
Figure 7: Overview of the tool state labeling process. Here the ToolState label is ”Not worn” or ”Worn”
The evaluation is based on multiple wear indicators, including workpiece surface quality, crater formation on the tool, chip shape and color, and audible changes in cutting noise. However, this labeling process is inherently subjective, as tool wear progression is complex and difficult to pinpoint with absolute temporal precision. Consequently, it is not possible to objectively determine the exact moment when wear begins to significantly degrade the milling process. This means that no direct accuracy or quality metric can be established to evaluate the data sets. Overall, the ToolState label is the most suitable option for wear prediction in this research.
@@ V Evaluation
It can be seen that there are only small differences between the training data of 80% and 60%. This indicates that a good prediction of machine wear can be achieved by training on 60% training data, which corresponds to three tool life cycles.
The analysis of model accuracy reveals that ConvNeXt-based models outperform classifiers that utilize manually designed features. It seems that the ConvNeXt backbone is more effective in extracting features that capture the wear behavior of the milling process. The ”DTreeC” and the ”SVC”-Model, may have limited representation capabilities, which restricts their ability to model complex wear patterns. Another key aspect influencing model performance is signal noise, particularly due to the sensor’s rotation. The CNN-ViT hybrid architecture handles this challenge more effectively by focusing more on spatial patterns rather than relying on exact values. Despite this, feature-based classifiers have significantly fewer parameters, making them computationally efficient while still delivering solid performance. Interestingly, the ConvNeXt model with a simple dense layer as a classifier outperforms its LSTM-enhanced counterpart. This suggests that the STFTs already contain sufficient temporal information, making additional sequence modeling unnecessary. Furthermore, short-term variations, such as sudden shocks from a cutting edge, appear to carry more predictive value than long-term trends, given that wear generally increases gradually. Another possible explanation is that LSTMs tend to overfit, especially when trained on a relatively small dataset, highlighting the risk of excessive model complexity without a proportional gain in performance.
Although the prediction accuracy or binary accuracy reflects the overall quality of the model, in reality it is more important how the model predicts wear over time, in particular when the model changes its prediction from not worn to worn. Fig. 9 shows the prediction of wear over time. The y-axis illustrates the categories of the binary classification, which are defined as follows:
$C:\{0,1\}\to\{\text{Not worn},\text{Worn}\}$ (10)
@@
The wear subjectively labeled by the operator (true label) is shown in blue and the model prediction is shown in green. Depending on the plot, several tool life cycles are shown concatenated in one plot. When the label changes from Worn to Not worn, a new tool life cycle begins. For post-processing, a simple filter is applied to the classification results to address the logical inconsistency that a tool cannot briefly appear worn and then revert to an unworn state. This filter eliminates a small proportion of such outliers from the plots, thereby enhancing both the accuracy over time and the readability of the visual representations. However, prominent outliers remain visible in the plots. Figure 9 illustrates that time accuracy improves as the size of the training set increases. It is noteworthy that certain models, such as ”ConvNeXt”, demonstrate relatively good time accuracy even with only 40% of the training data. The ”SVC”, despite its simplicity, achieves only slightly lower time accuracy compared to more advanced models. The results together demonstrate that a minimum of two, preferably three, tool life cycles is required to train a model with the capacity to make reliable predictions. This assumption is based on the premise that the transition from no wear to worn occurs with minimal time deviations. In practice, recognizing wear too late is undesirable. To mitigate this, the transition from ”not worn” to ”worn” could be adjusted slightly earlier in the labeling process, thereby introducing a calculated bias for the early prediction of wear. Alternatively, if avoiding such a bias is preferred and high time accuracy is still desired, training on data from four or more tool life cycles becomes necessary.
@@
It should be noted that the training data of the different splits are identical to those used in the heatmap in Fig. 8. It is evident that the models on the unseen machine exhibit reduced performance, yet the ConvNeXt model continues to demonstrate the most optimal outcomes. Interestingly, models trained with the largest dataset (80% training data) perform worse than those trained with smaller datasets. This suggests that the change of a machine has fundamentally altered the wear behavior, causing models trained on extensive data to overfit to the specific characteristics of the original machine. Models trained with 40% of the training data achieve the highest accuracy. An examination of the accuracy over time (Fig. 11) reveals that the models are capable of detecting changes in the DMU60FD dataset and successfully predict wear in some tool life cycles.
@@
Nevertheless, when considering all models collectively, while each tool life cycle is successfully recognized by at least one model, no single model achieves consistent accuracy across all cycles. This suggests that the findings derived from the training dataset can be applied to the DMU60FD dataset. However, no overarching patterns are identified that apply to all tool life cycles. This discrepancy may be attributed to the differing wear behavior in the frequency domain between the two machines, as illustrated in Fig. 12. For the Chiron, significant changes occur primarily below 4000 Hz, whereas in the DMU60FD, wear-related variations are also observed above 4000 Hz. It is possible that the models focus predominantly on the sub-4000 Hz range, where the DMU60FD exhibits only partial changes. Consequently, the models may encounter difficulties in fully capturing wear patterns on the DMU60FD, as the most relevant wear-related information for this machine may lie above 4000 Hz.
Figure 12: Display of two FFTs with a worn tool, each from the Chiron and DMU60FD
In summary, the fundamental issue associated with TCM’s is also clear in this study, namely that the complex characteristics of each machine make it difficult to implement one prediction model across other unseen machines. To improve predictions, the models require additional training data from the specific machine.
@@ VI Conclusion
This paper addresses the question of how to develop a TCM system that can be transferred to industrial applications. For this, a modular approach and a non-invasive setup is employed so that a TCM can be installed for different milling processes with minimal effort. For a non-invasive setup, a 1D acceleration sensor is used, and for modality the sensor is integrated into a toolholder. In this research, the acceleration signal is processed in both the time domain and the frequency domain, to generate valuable data input for wear prediction or TCM. The best performing TCM approach uses short-time Fourier transform as input and automatic feature extraction through a transformer-inspired CNN architecture to predict wear. Given that the development of a TCM is more cost-effective when fewer training data are required, the study also investigates the amount of training data needed from a milling process to get a sufficient accuracy: To achieve the highest degree of accuracy in wear prediction, at least four tool life cycles should be included in the training data, two or three cycles for an acceptable accuracy. Among the models tested, the ConvNeXt model delivered the best results with 99% accuracy. Additionally, the models capacity to predict wear for an untrained milling machine is evaluated. Although models demonstrate some ability to transfer information and predict wear for certain tools, they struggle to fully capture the inherent complexity and variability of new milling machines, resulting in inconsistent wear predictions. This underscores the necessity for transfer learning to adapt the model to different machining conditions. Concurrent with the publication of this article, ongoing research is investigating the amount of data required for effective transfer learning, determining which models adapt best, and refining training weight adjustments. Additionally, efforts are underway to explore efficient methods for industrial data collection and labeling to facilitate real-world deployment.
@@ Appendix A Tables of this research
Mean $\frac{1}{N}\sum_{i=1}^{N}x_{i}$, remains stable in an undisturbed process but shifts with wear [9].
RMS (Root Mean Square) $\sqrt{\frac{1}{N}\sum_{i=1}^{N}x_{i}^{2}}$, measures signal energy, higher values indicate increased wear [41].
Standard Deviation (SD) $\sqrt{\frac{1}{N-1}\sum_{i=1}^{N}(x_{i}-\bar{x})^{2}}$, quantifies signal variation and fluctuation intensity [42].
Crest Factor $\frac{\max|x|}{\text{RMS}}$, detects impact forces and transient events in the signal [43].
Kurtosis $\frac{1}{N}\sum_{i=1}^{N}\left(\frac{x_{i}-\bar{x}}{\text{SD}}\right)^{4}$, measures impulsiveness and sharp peaks [44].
Skewness $\frac{1}{N}\sum_{i=1}^{N}\left(\frac{x_{i}-\bar{x}}{\text{SD}}\right)^{3}$, indicates asymmetry in the signal distribution [45].
Statistical Mode $\text{Mode}(X)=\arg\max_{x\in X}f_{cnt}(x)$, identifies the most frequent value, robust to outliers.
SD of Statistical Mode Measures the variability of the mode, indicating process stability.
Peak-to-Peak (P2P) $\max(x)-\min(x)$, tracks extreme signal values, sensitive to transient events [46].
Center Frequency $\frac{\sum f_{i}*A_{i}}{\sum A_{i}}$, average of the spectral content, shows dynamic changes [47]
Dominant Frequency $\arg\max_{f}(PSD(f))$, the frequency with the highest energy, can shift due to wear [48]
TABLE I: Statistical Features and corresponding formulas. The sources refer to research papers that have used a comparable feature.
Feature Name Formula / Description
Energy of Power Spectral Density (PSD) $\frac{\sum{|X(f)|^{2}}}{N}$, where $X(f)$ is the PSD of $x(t)$, reflects vibration energy and wear progression [49].
Power Spectral Entropy $-\sum_{n=1}^{N}P(n)\log_{2}P(n)$, where $P(n)$ is the probability distribution, indicating frequency spread due to wear [50].
Periodic Energy $\sum_{f\in F_{\text{process}}}E_{f}$, energy in harmonic components (e.g., spindle, tooth-pass frequencies) [51].
Aperiodic Energy $E_{\text{total}}-\sum_{f\in F_{\text{process}}}E_{f}$, quantifies energy from non-harmonic, ”chaotic” components (e.g., friction).
Relative Aperiodic Energy $\frac{E_{\text{ap}}}{E_{\text{total}}}$, ratio of aperiodic energy to total energy, indicating noise and disturbances.
Auto-Correlation $\frac{\sum_{i=1}^{N-\tau}(x_{i}-\bar{x})(x_{i+\tau}-\bar{x})}{\sum_{i=1}^{N}(x_{i}-\bar{x})^{2}}$, signal randomness and periodicity [52].
Higuchi Fractal Dimension Estimated via linear regression on $\{(\log(1/k),\log(L(k)))\}$, where $k$ is the scale parameter and $L(k)$ the curve length. Detects time series complexity [53].
TABLE II: Description of advanced features, mostly for the frequency domain. The sources refer to research papers that have used a comparable feature.
@@
TABLE III: Process parameters from the measured acceleration data for the models